1. If 22.0 grams of propane (C3H8) is combusted in a bomb calorimeter containing 3.48 liters of water, calculate the enthalpy (ΔH) of propane if the water temperature rises 26.5 °C?f water to create a 3.5% by mass solution?

2. What mass of NaF must be mixed with 25 mL of water to create a 3.5% by mass solution?
Answer #1

heat gained by water(q) = m*s*DT

m = mass of water = d*v = 3480*1 = 3480 g

s = specific heat of water = 4.184 j/g.c

DT = 26.5

q = 3480*4.184*26.5 = 385.85 kj

No of mol of propane burned = 22.0/44 = 0.5 mol

DHcombustion of propane = -q/n = -385.85/0.5 = -771.7 kj/mol

2)
%(m/m) means , mass of solute present in 100 g solution.

mass percent of solution(m/m) = mass of solute/mass of solution*100

mass of solute(NaF) = x g

mass of solution = NaF + water

           = x + 25      ( density of water = 1 g/ml)

3.5 = (x/(25+x))*100

x = 0.91

mass of solute(NaF) = x = 0.91 g